# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ 5 Experiments
We evaluate our method, STAD, against various baselines on a range of datasets under challenging settings. In Sec. 5.1, we study temporal distribution shifts as defined in Sec. 2, demonstrating the difficulty of the task and STAD’s robustness in practical settings. In Sec. 5.2, we go beyond temporal shifts and find that STAD is competitive on reproduction datasets and synthetic corruptions as well. Finally, in Sec. 5.3, we provide insights into STAD’s mechanisms, confirming the reliability of its prototypes and highlighting the importance of modeling shift dynamics through an ablation study. We now describe the datasets, source architectures, and baselines. Further implementation details can be found in App. C.
@@
Our primary focus is on temporal distribution shifts. To further evaluate the effectiveness of our method, we also test its performance on reproduction datasets and standard image corruptions.
  • EVIS: the evolving image search (EVIS) dataset (Zhou et al., 2022) consists of images of 10 electronic product and vehicle categories retrieved from Google search, indexed by upload date. The dataset captures shift caused by rapid technological advancements, leading to evolving designs across time. It includes 57,600 RGB images of 256x256 pixels from 2009 to 2020. Models are trained on images from 2009-2011 and evaluated on images from 2012-2020.
@@
We employ a variety of source architectures to demonstrate the model-agnostic nature of our method. They notably vary in backbone architecture (we use CNN, DenseNet, ResNet, WideResNet) and dimensionality of the representation space (from 32 up to 2048). We list details in Sec. C.1. In addition to the source model, we compare against 7 baselines representing fundamental approaches to TTA. Five of them adapt the feature extractor: Batch norm adaptation (BN) (Schneider et al., 2020; Nado et al., 2020), TENT (Wang et al., 2021), CoTTA (Wang et al., 2022), SHOT (Liang et al., 2020) and SAR (Niu et al., 2023). Like our method STAD, two baselines adapt the last linear layer: T3A (Iwasawa & Matsuo, 2021) and LAME (Boudiaf et al., 2022). More details are provided in Sec. C.2. Batch sizes are the same for all baselines (Sec. C.3). To ensure optimal performance on newly studied datasets, we conduct an extensive hyperparameter search for each baseline, following the hyperparameters and value ranges suggested in the original papers. We then report the best settings. Sec. C.3 specifies the grid searches we conducted.
@@ 5.1 Temporal Distribution Shifts
We start by evaluating the adaptation abilities to temporal distribution shift on three image classification datasets (Yearbook, EVIS, FMoW-Time), which vary in number of classes (2, 10, 62, respectively), representation dimension (32, 512, 1024, respectively) and shift dynamics (recurring, progressive and rapid, respectively as visible in Fig. 4). For the low-dimensional representations of Yearbook, we also evaluate our computationally costly Gaussian model (STAD-Gauss). We evaluate two settings: (i) covariate shift with a uniform label distribution and (ii) covariate shift with additional shift in the label distribution $\mathbb{Q}_{t}({\mathbf{y}})$. Having a uniform label distribution—samples are evenly shuffled, making test batches nearly class balanced—has been the standard evaluation setting for TTA. However, particularly in temporal distribution shifts, it is highly unlikely that samples arrive in this iid-manner. Instead, temporally correlated test streams often observe consecutive samples from the same class (Gong et al., 2022). We follow Lim et al. (2023), ordering the samples by class and thus inducing an extreme label shift. We draw the class order uniformly at random.
Table 2: Accuracy on temporal distribution shifts and label shifts, averaged over three random seeds. Colors highlight performance gains and drops relative to the source model. Best model in bold, second-best underlined.
@@
Figure 4: Accuracy over time for three temporal distribution shifts: On Yearbook for example, STAD can mitigate the distribution shift by improving 10 points upon the source model for some years (1980s). Some baselines perform similarly, which is depicted by overlaying accuracy trajectories. Error bars for FMoW-Time are tiny.
@@
Tab. 2 shows overall accuracy, averaged over all time steps and three random training seeds. Results that do not outperform the source model are highlighted in red and ones that do in blue. Methods that primarily adapt the feature extractor are shown in the upper section of the table. Ones that, like ours, adapt the classifier are shown in the lower section. To summarize the results: on Yearbook, all methods perform well without label shift, and with label shift, only classifier-based methods improve upon the source baseline. Feature-based methods completely fail on EVIS, and all models, except LAME and STAD-vMF under label shift, fail on FMoW-Time. This leads us to three key takeaways: first, these TempTTA tasks are inherently difficult, leading to smaller adaptation gains overall compared to traditional corruption experiments. Second, methods that adapt only the last layer clearly perform better on temporal distribution shifts under both label distribution settings. This indicates that perhaps ‘less is more’ for TempTTA. Third, STAD demonstrates the most consistent performance, ranking as the best or second-best model across all datasets and settings. On Yearbook, both the Gaussian and vMF variants outperform the baselines, with the fully parameterized Gaussian model better capturing the distribution shift than the more lightweight vMF model. Fig. 4 displays adaptation performance over different timestamps. We see that on EVIS (middle) the methods markedly separate, which reflects the aforementioned gap between feature-based and classifier-based approaches. The reader may also wonder if STAD can be stacked on top of a feature-based approach. We present results exemplary for BN in Tab. 8 (Sec. D.2) but found that it offers no significant improvement to STAD’s performance.
@@
Tab. 2 demonstrates that STAD performs particularly well under imbalanced label distributions, delivering the best results on both Yearbook and FMoW-Time. This advantage stems from STAD’s clustering approach, where a higher number of samples from the same ground truth class provides a stronger learning signal, leading to more accurate prototype estimates. This is particularly notable on FMoW, where STAD improves upon the source model by more than 17 points. Further, the performance gap between classifier and feature extractor adaptation methods becomes even more pronounced in this setting. This is not surprising as the latter typically depend heavily on current test-batch statistics, making them vulnerable to imbalanced class distributions (Niu et al., 2023). In contrast, having fewer classes to cause confusion allows classifier-based methods to benefit from label shift, with STAD delivering the most persistent adaptation gains.
Figure 5: Batch size effects under covariate shift (first row) and additional label shift (second row): Compared to most baselines, STAD-vMF (dark blue) is relatively robust to very small batch sizes. For label shift on EVIS and FMoW-Time, we observe a sweet spot around batch size 16.
@@
Adapting to a small number of samples is crucially valuable, as one does not have to wait for a large batch to accumulate in order to make adapted predictions. We next evaluate performance across 12 different batch sizes ranging from 1 to 2048 under both covariate shift and additional label shift. Fig. 5, displays results. STAD-vMF (dark blue line) is able to maintain stable performance under all batch sizes. In Tab. 7 (Sec. D.1), we report values for batch size 1 showing that STAD adapts successfully even in the most difficult setting. In contrast, methods relying on normalization statistics collapse when not seeing enough samples per adaptation step. For example, on FMoW, feature-based methods collapse to nearly random guessing at the smallest batch sizes. When batch sizes are large, note that TENT, CoTTA, SHOT, and SAR hit memory constraints on FMoW-Time, failing to close the gap to the source model performance in the label shift scenario.
@@ 5.2 Beyond Temporal Shifts: Reproduction Datasets and Synthetic Corruptions
Table 3: Accuracy on reproduction datasets and label shifts, averaged over three random data seeds.
@@
We evaluate our method on reproduction datasets (CIFAR-10.1 and ImageNetV2), which have recently gained attention as benchmarks for more realistic and challenging distribution shifts (Zhao et al., 2023b). We use a batch size of 100 for CIFAR-10.1 and 64 for ImageNetV2. Tab. 3 confirms the difficulty of adapting to more natural distribution shifts. For CIFAR-10.1, only T3A and STAD outperform the source model for both with and without label shift, with STAD adapting best. For ImageNetV2, none of the methods improve upon the source model when the label distribution is uniform. We again observe that classifier-adaptation methods handle label shifts better by a significant margin.
Table 4: Accuracy on CIFAR-10-C across different levels of synthetic corruptions
@@
Lastly, we test our method on gradually increasing noise corruptions of CIFAR-10-C, a standard TTA benchmark. We use a batch size of 100. Tab. 4 shows the accuracy averaged across all corruption types. We make three key observations. First, performance gains are much higher than on previous datasets indicating the challenge posed by non-synthetic shifts. Second, as expected, methods adapting the backbone model are more performative on input-level noise, since such shifts primarily affect earlier layers (Tang et al., 2023; Lee et al., 2023). Lastly, STAD is consistently the best method amongst those adapting only the last linear layer.
@@ 5.3 Analysis of Tracking Abilities
We evaluate how well STAD’s inferred cluster centers align with the ground truth cluster centers (computed using labels). We chose the progressively increasing distribution shift of CIFAR-10-C as this dataset represents a bigger challenge for STAD. Fig. 6 (left) shows distance (in angular degrees) to the ground truth cluster centers for both the source model and STAD. STAD (blue line) adapts effectively, significantly reducing the angular distance to the ground truth cluster centers. For the source model, the progressive distribution shift causes the ground truth cluster centers to drift increasingly further from the source prototypes (grey line). Additionally, by computing dispersion (Ming et al., 2023) (Fig. 6, middle), which measures the spread of the prototypes (in angular degrees), we find that STAD mirrors the ground truth trend (yellow line) of clusters becoming closer together. This is a promising insight, as it suggests that STAD’s cluster dispersion could potentially serve as an unsupervised metric to proactively flag when clusters start overlapping and estimate adaptation accuracy. In Fig. 6 (right), we plot accuracy vs dispersion of STAD’s prototypes for different corruptions and severity levels, confirming that they positively correlate.
@@
STAD is proposed with the assumption that adapting the class prototypes based on those of the previous time step facilitates both rapid and reliable adaptation. However, one could also consider a static version of STAD that does not have a transition model (Eqn. 1). Rather, the class prototypes are computed as a standard mixture model (Eqn. 2) and without considering previously inferred prototypes. Tab. 5 presents the accuracy differences between the static and dynamic versions of STAD in percentage points. Removing STAD’s transition model results in a substantial performance drop of up to 28 points. This supports our assumption that state-space models are well-suited for the task of TempTTA.
6 Conclusion
We have presented State-space Test-time ADaptation (STAD), a novel test-time adaptation strategy based on probabilistic state-space models that addresses the challenges of temporal distribution shifts. Our Gaussian and vMF variants of STAD effectively track the evolution of the last layer under distribution shifts, enabling unsupervised adaptation in deployed models. Our extensive experiments highlight the significant challenges that temporal distribution shifts present for existing TTA methods. Future work on TempTTA could explore incorporating temporal information, such as timestamps, to better model the passage of time and extend the approach to irregularly sampled settings.
Acknowledgments
We thank Metod Jazbec and Rajeev Verma for helpful discussions and feedback on the draft. This project was generously supported by the Bosch Center for Artificial Intelligence. No Johns Hopkins University resources were used to conduct this research.
@@ Acknowledgments
  • App. A provides details on the Gaussian formulation of STAD (STAD-Gauss) stating the employed transition and emission model.
@@
  • App. B details the von Mises-Fisher formulation (STAD-vMF). In Sec. B.1, we provide the inference scheme including all update equations for the variational EM step. In Sec. B.2, we provide the algorithmic overview of STAD-vMF.
  • App. C contains various implementation details.
  • App. D shows additional experimental results on single sample adaptation (Sec. D.1), STAD in combination with BN (Sec. D.2), visualizations of representation space (Sec. D.3) runtime comparisons (Sec. D.4) and sensitivity to hyperparameters (Sec. D.5).
Appendix A Details on STAD-Gaussian
A.1 Model Formulation

### v3

@@ 5 Experiments
We evaluate our method, STAD, against various baselines on 7 datasets under challenging settings. Sec.˜5.1 studies temporal distribution shifts as defined in Sec.˜2, demonstrating the difficulty of the task and STAD’s robustness in practical settings. In Sec.˜5.2, we go beyond temporal shifts and find that STAD is competitive on reproduction datasets and synthetic corruptions as well (see Sec.˜E.1 for results on domain adaptation). Finally, in Sec.˜5.3, we provide insights into STAD’s mechanisms, confirming the reliability of its prototypes and highlighting the importance of modeling shift dynamics through an ablation study. We now describe the experimental setup. Details are listed in App.˜D.
@@
We investigate temporal distribution shifts using three image classification datasets spanning several years. Yearbook (Ginosar et al., 2015) involves binary gender prediction on portrait images, capturing changes in demographics and beauty standards over time. EVIS (Zhou et al., 2022a) categorizes vehicles and electronic products. FMoW-Time (Koh et al., 2021) maps satellite images to land use categories. Each dataset comprises samples from multiple years, with the earlier years used for training and the later years for testing (see Sec.˜D.1 for details). To evaluate the effectiveness of our method beyond TempTTA, we also test its performance on reproduction datasets (CIFAR-10.1, ImageNetV2) and image corruptions (CIFAR-10-C).
@@
We employ a variety of source architectures to demonstrate the model-agnostic nature of STAD. They vary in backbone architecture (ViT, CNN, DenseNet, ResNet, WideResNet) and dimensionality of the representation space (from 32 up to 2048). We list details in Sec.˜D.2. We compare against 8 TTA baselines representing fundamental TTA approaches. Six of them adapt the feature extractor: Batch norm adaptation (BN) (Schneider et al., 2020; Nado et al., 2020), TENT (Wang et al., 2021), CoTTA (Wang et al., 2022b), SHOT (Liang et al., 2020) SAR (Niu et al., 2023) and RoTTA (Yuan et al., 2023). Like our method STAD, two baselines adapt the last layer: T3A (Iwasawa & Matsuo, 2021) and LAME (Boudiaf et al., 2022). More details are provided in Sec.˜D.3. Batch sizes are the same for all baselines. To ensure optimal performance on newly studied datasets, we conduct an extensive hyperparameter search for each baseline (see Sec.˜D.4) and report the best setting.
@@ 5.1 Temporal Distribution Shifts
Table 2: Accuracy on temporal distribution shifts and label shifts, averaged over three random seeds. Colors highlight performance that either improves or degrades relative to the source model. Best model in bold, second-best underlined.
@@
We start by evaluating the adaptation abilities to temporal distribution shift on three image classification datasets (Yearbook, EVIS, FMoW-Time), which vary in number of classes (2, 10, 62, respectively), representation dimension (32, 512, 1024, respectively) and shift dynamics (recurring, progressive and rapid, respectively as visible in Fig.˜3). For the low-dimensional representations of Yearbook, we also evaluate our computationally costly Gaussian model (STAD-Gauss). We evaluate two settings: (i) covariate shift with a uniform label distribution and (ii) covariate shift with additional shift in the label distribution $\mathbb{Q}_{t}({\mathbf{y}})$. Having a uniform label distribution—samples are evenly shuffled, making test batches nearly class balanced—has been the traditional evaluation setting for TTA. However, particularly in temporal distribution shifts, it is highly unlikely that samples arrive in this iid-manner. Instead, temporally correlated test streams often observe consecutive samples from the same class (Gong et al., 2022). We follow Lim et al. (2023), ordering the samples by class and thus inducing an extreme label shift. We draw the class order uniformly at random.
@@
Tab.˜2 shows overall accuracy, averaged over all time steps and three random training seeds. Results that do not outperform the source model are highlighted in red and ones that do in blue. Methods that primarily adapt the feature extractor are shown in the upper section of the table. Ones that, like ours, adapt the classifier are shown in the lower section. To summarize the results: on Yearbook, all methods perform well without label shift, and with label shift, only classifier-based methods improve upon the source baseline. Feature-based methods completely fail on EVIS, and all models, except LAME and STAD-vMF under label shift, fail on FMoW-Time. This leads us to three key takeaways: first, these TempTTA tasks are inherently difficult, leading to smaller adaptation gains overall compared to traditional corruption experiments. Second, methods that adapt only the last layer clearly perform better on temporal distribution shifts under both label distribution settings. This indicates that perhaps ‘less is more’ for TempTTA. Third, STAD demonstrates the most consistent performance, ranking as the best or second-best model across all datasets and settings. On Yearbook, both the Gaussian and vMF variants outperform the baselines, with the fully parameterized Gaussian model better capturing the distribution shift than the more lightweight vMF model. Fig.˜3 displays adaptation performance over different timestamps. We see that on EVIS (middle) the methods markedly separate, which reflects the aforementioned gap between feature-based and classifier-based approaches. The reader may wonder if STAD can be stacked on a feature-based approach. We present results for BN in Tab.˜11 (Sec.˜E.3) but find no significant improvement in STAD’s performance.
Figure 3: Accuracy over time for TempTTA: STAD mitigates distribution shifts by improving up to 10 points over the source model (Yearbook, 1980s). Some baselines perform similarly, shown by overlaying accuracy trajectories.
@@
Tab.˜2 demonstrates that STAD performs particularly well under imbalanced label distributions, delivering the best results on both Yearbook and FMoW-Time. This advantage stems from STAD’s clustering approach, where a higher number of samples from the same ground truth class provides a stronger learning signal, leading to more accurate prototype estimates. This is particularly notable on FMoW, where STAD improves upon the source model by more than 17 points. Further, the performance gap between classifier and feature extractor adaptation methods becomes even more pronounced in this setting. This is not surprising as the latter typically depend heavily on current test-batch statistics, making them vulnerable to imbalanced class distributions (Niu et al., 2023). In contrast, having fewer classes to cause confusion allows classifier-based methods to benefit from label shift, with STAD delivering the most persistent adaptation gains.
Figure 4: Batch size effects under covariate shift (first row) and additional label shift (second row): STAD-vMF (dark blue) shows robustness to small batches, with a sweet spot around batch size 16 for label shift on EVIS and FMoW-Time.
@@
Adapting to a small number of samples is crucially valuable, as one does not have to wait for a large batch to accumulate in order to make adapted predictions. We next evaluate performance across 12 different batch sizes ranging from 1 to 2048 under both covariate shift and additional label shift. Fig.˜4, displays results. STAD-vMF (dark blue) is able to maintain stable performance under all batch sizes. In Tab.˜10 (Sec.˜E.2), we report values for batch size 1 showing that STAD adapts successfully even in the most difficult setting. In contrast, methods relying on normalization statistics collapse when not seeing enough samples per adaptation step. For example, on FMoW, feature-based methods collapse to nearly random guessing at the smallest batch sizes. When batch sizes are large, note that TENT, CoTTA, SHOT, and SAR hit memory constraints on FMoW-Time, failing to close the gap to the source model performance in the label shift scenario.
@@ 5.2 Beyond Temporal Shifts: Reproduction Datasets and Synthetic Corruptions
Table 3: Accuracy on reproduction datasets.
@@
We evaluate our method on reproduction datasets (CIFAR-10.1, ImageNetV2), which have been considered as more realistic and challenging distribution shifts (Zhao et al., 2023b). Tab.˜3 confirms the difficulty of adapting to more natural distribution shifts. For CIFAR-10.1, only T3A and STAD outperform the source model for both with and without label shift, with STAD adapting best. For ImageNetV2, none of the methods improve upon the source model when the label distribution is uniform. We again observe that classifier-adaptation methods handle label shifts better by a significant margin.
Table 4: Accuracy on synthetic corruptions (CIFAR-10-C)
@@
Lastly, we test our method on gradually increasing noise corruptions of CIFAR-10-C, a standard TTA benchmark and provide results on ImageNet-C in Sec.˜E.4. Tab.˜4 shows the accuracy averaged across all corruption types for CIFAR-10-C. We make three key observations. First, performance gains are much higher than on previous datasets indicating the challenge posed by non-synthetic shifts. Second, as expected, methods adapting the backbone model are more performative on input-level noise, since such shifts primarily affect earlier layers (Tang et al., 2023; Lee et al., 2023). Lastly, STAD is consistently the best method amongst those adapting only the last layer.
@@ 5.3 Analysis of Tracking Abilities
We evaluate how well STAD’s inferred cluster centers align with the ground truth cluster centers (computed using labels). We chose the progressively increasing distribution shift of CIFAR-10-C as this dataset represents a bigger challenge for STAD. Fig.˜5 (left) shows distance (in angular degrees) to the ground truth cluster centers for both the source model and STAD. STAD (blue line) adapts effectively, significantly reducing the angular distance to the ground truth cluster centers. For the source model, the progressive distribution shift causes the ground truth cluster centers to drift increasingly further from the source prototypes (red line). Additionally, by computing dispersion (Ming et al., 2023) (Fig.˜5, middle), which measures the spread of the prototypes (in angular degrees), we find that STAD mirrors the ground truth trend (yellow line) of clusters becoming closer together. This is a promising insight, as it suggests that STAD’s cluster dispersion could potentially serve as an unsupervised metric to proactively flag when clusters start overlapping and estimate adaptation accuracy. In Fig.˜5 (right), we plot accuracy vs dispersion of STAD’s prototypes for different corruptions and severity levels, confirming that they positively correlate.
@@
STAD is proposed with the assumption that adapting the class prototypes based on those of the previous time step facilitates rapid and reliable adaptation. However, one could also consider a static version of STAD that does not have a transition model (Eqn.˜1). Rather, the class prototypes are computed as a standard mixture model (Eqn.˜2) without considering previously inferred prototypes. Tab.˜5 presents the accuracy differences between the static and dynamic versions of STAD. Removing STAD’s transition model results in a substantial performance drop of up to 28 percentage points. This supports our assumption that SSMs are well-suited for TempTTA.
@@
6 Discussion and Conclusion
We studied temporal distribution shifts and demonstrated the significant challenges they pose for existing TTA methods. We proposed STAD, a novel TTA strategy based on probabilistic state-space models to address temporal shifts. Our Gaussian and vMF variants of STAD effectively track the evolution of the last linear layer under distribution shifts, enabling unsupervised adaptation in deployed models. We found STAD to be most effective for structural temporal shifts and label shifts (Tabs.˜2 and 3). While effective across a range of settings, STAD’s design inherits limitations regarding the types of distribution shift it can mitigate. Notably, its performance depends on the shift being visible in the last layer, as adapting only the final classifier is less effective when earlier layers are primarily affected—a behavior we observe for synthetic corruptions in CIFAR-10-C (Table 4) and ImageNet-C (Table 12), consistent with prior work (Lee et al., 2023). STAD also assumes that last-layer representations change gradually over time, and that class prototypes remain temporally correlated. This makes abrupt shifts more challenging to address, though in our experiments we did not encounter a scenario where sudden shifts fully broke the temporal correlation (see Sec.˜E.1). Future work on TempTTA could incorporate timestamps to better model time progression or determine when adaptation is no longer feasible.
Broader Impact Statement
This work introduces a probabilistic state-space framework for TTA under temporal distribution shifts. By modeling time-evolving feature dynamics and inferring class prototypes without labels, STAD enables continuous model adaptation in settings where labels are unavailable. STAD contributes to more robust and flexible deployment of machine learning systems in non-stationary environments. However, because the method operates without ground-truth labels, there is a risk of compounding errors if the inferred prototypes drift too far from reality—especially in cases of abrupt or adversarial shifts. Overreliance on unsupervised adaptation without safeguards could lead to silent model degradation. Practitioners should therefore monitor model behavior carefully and understand the assumptions underlying the temporal dynamics. Overall, this work supports the development of adaptive models under distribution shift.
Acknowledgment
We thank Metod Jazbec for valuable discussions and helpful feedback on the draft, and we thank Rajeev Verma for constructive comments on an early manuscript. This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not utilize resources from Johns Hopkins University for this project.
Appendix A Appendix
@@
  • App.˜B discusses further related work such as adjacent robustness paradigms and further TTA settings that aim for realistic evaluations.
  • App.˜C provides methodological details on STAD.
    • Sec.˜C.1 defines the Gaussian SSM (STAD-Gauss), provides the respective EM update equations and an algorithmic overview (Algorithm˜2).
@@
    • Sec.˜C.2 details the von Mises-Fisher SSM (STAD-vMF). Here, we notably provide the inference scheme including all update equations for the variational EM step as well as an algorithmic summary (Algorithm˜3).
  • App.˜D contains various implementation details.
  • App.˜E provides additional experimental results on
Appendix B Expanded Related Work
Table 6: Comparison of realistic TTA settings and associated methods. We distinguish settings with respect to assumptions on the feature distribution and label distribution as well as if inherently time-ordered data streams and single sample settings have been considered. Notably, we categorize the assumption on the feature distribution into stationary (a single, static feature distribution), non-stationary (continually evolving distribution) and mixture (mixture distribution, e.g. samples of different corruption types observed simultaneously).
TTA setting Related method Feature distribution $\mathbb{Q}({\mathbf{x}})$ Label distribution $\mathbb{Q}({\textnormal{y}})$ Time-ordered Single sample
Fully TTA TENT (Wang et al., 2021) stationary balanced
Continual TTA CoTTA (Wang et al., 2022b) non-stationary balanced
Non-iid TTA LAME (Boudiaf et al., 2022) stationary imbalanced
NOTE (Gong et al., 2022) stationary imbalanced
Practical TTA RoTTA (Yuan et al., 2023) non-stationary imbalanced
Wild TTA SAR (Niu et al., 2023) mixture imbalanced
N/A TTN (Lim et al., 2023) stationary, non-stationary, mixture imbalanced
N/A RMT (Döbler et al., 2023) non-stationary balanced
Universal TTA ROID (Marsden et al., 2024) non-stationary, mixture imbalanced
Real-world TTA TRIBE (Su et al., 2024) non-stationary imbalanced
UniTTA UniTTA (Du et al., 2025) non-stationary, mixture imbalanced
TempTTA STAD (ours) non-stationary imbalanced
Realistic TTA
aims to evaluate models under challenging test conditions that may occur in real-world deployments. We provide a tabular overview of prominent and recent work in Tab.˜6. For a more comprehensive review of TTA settings and methods, see Xiao & Snoek (2024). Since TTA methods typically rely on test data statistics, they are often sensitive to the ordering of samples in the test data stream. As a result, prior work has explored the robustness of TTA under different sample ordering strategies. Two key factors influence sample order: the domain index, which determines the feature distribution $\mathbb{Q}({\mathbf{x}})$, and the class index, which determines the label distribution $\mathbb{Q}({\textnormal{y}})$.
In the standard fully TTA setting (Wang et al., 2021), the test stream consists of a single domain, and test data is sampled i.i.d., resulting in a uniform label distribution per batch. Continual TTA (Wang et al., 2022b) extends this setting by considering multiple domains sequentially, leading to a non-stationary feature distribution. Non-i.i.d. TTA (Boudiaf et al., 2022; Gong et al., 2022; Zanella et al., 2025) is another extension of fully TTA that challenges the i.i.d. assumption by introducing temporal correlation in the sampling procedure, causing an imbalanced label distribution per test batch. The Practical TTA setting (Yuan et al., 2023) combines continual TTA and non-i.i.d. TTA by incorporating both continually changing domains and temporal correlation. Wild TTA (Niu et al., 2023) is another extension of non-i.i.d. TTA that additionally studies mixed domains (see also Universal TTA (Marsden et al., 2024)) and single sample adaptation (see also Lim et al. (2023); Döbler et al. (2023); Marsden et al. (2024)). Real-world TTA (Su et al., 2024) builds on practical TTA by also controlling the global label distribution across the data stream. Lastly, UniTTA (Du et al., 2025) comprises 36 sampling strategies, considering ordering and imbalance of both domains and class labels.
While past work has studied adaptation to non-stationary distributions, they mostly consider a sequence of categorical domains (e.g. different corruption types). In contrast, TempTTA focuses on sequences where the domain index is temporal. The data stream thus follows an inherent time-ordering (Bai et al., 2023). To test the applicablity of our method in challenging settings, we also evaluate single sample adaptation and class-imbalanced label distributions in Sec.˜5.1.
Domain Generalization (DG)
The goal of DG (Zhou et al., 2022a) is to learn a predictive model that can generalize well to any unseen domains, assuming access to multiple source domains at the training stage. Examples of common approaches are domain-invariant feature learning (Arjovsky et al., 2019), data augmentation (Zhang, 2018) and regularization (Krueger et al., 2021). The most relevant subfield to this work is temporal DG (TDG) (Nasery et al., 2021; Qin et al., 2022; Bai et al., 2023; Xie et al., 2024b; Jin et al., 2024; Zeng et al., 2024; Cai et al., 2024; Xie et al., 2024a), which models dynamics from sequential source domains to generalize to evolving target domains. While both TDG and TempTTA address temporal shifts, TDG operates during training and requires a sequence of labeled source domains, whereas our approach improves the performance of arbitrary pre-trained models at test time using only a stream of unlabeled data.
Unsupervised Domain Adaptation (UDA)
UDA improves generalization by exploiting both labeled source data and unlabeled target data. Most related to our setting is gradual domain adaptation (GDA) (Hoffman et al., 2014; Wulfmeier et al., 2018; Bobu et al., 2018; Kumar et al., 2020; Abnar et al., 2021; Wang et al., 2022a), which aims to adapt to a target domain by exploiting intermediate domains with gradual distribution shift between source and target. GDA methods rely on access to both source and target data for distribution alignment. However, in many practical scenarios, source data may be unavailable due to e.g. privacy concerns, motivating the need for TTA.
Appendix C Methodological Details
In the following, we provide details on the probabilistic model and inference of STAD. Fig.˜6 displays the plate diagram.
${\mathbf{c}}_{t-1,n}$${\mathbf{c}}_{t,n}$${\mathbf{h}}_{t-1,n}$${\mathbf{h}}_{t,n}$${\mathbf{w}}_{t-1,k}$${\mathbf{w}}_{t,k}\ldots\ldots$$\pi_{t-1,k}$$\pi_{t,k}$$\psi^{\mathrm{trans}}$$\psi^{\mathrm{ems}}$ $N_{1}$ $N_{2}$ $K$
C.1 Details on STAD-Gaussian
Gaussian State-Space Model